A positive integer is called a palindrome if it reads the same forward as it does backward. For example, 959 and 8228 are palindromes , whereas 1332 is not neither the first nor the last of a palindrome can be 0
1)
Which of the following integers is a palindrome?
(A) 550 (B) 2255 (C) 2525 (D) 2552 (E) 5002
2)
How many three-digit palindromes are there ?
(A) 19 (B) 20 (C) 90 (D) 100 (E) 810
3)
The next two palindromes greater than 50805 are m and p and m < p . what is the value of p – m
(A) 10 (B) 90 (C) 100 (D) 110 (E) 210

Answer by Edwin McCravy(20086)


A positive integer  is called a palindrome if it reads the same forward as it
does backward. For example,  959 and 8228 are palindromes , whereas 1332 is
not  neither the first nor the last of a palindrome  can be 0
1)
Which of the following integers is a palindrome?
(A)   550   (B) 2255  (C) 2525   (D) 2552  (E)  5002
2)

(D) 2552 reads the same left to right as it reads right to left.



How many three-digit palindromes are there ?
(A)  19  (B) 20  (C)  90 (D) 100   (E) 810
3)

They have the form XYX

There are 9 choices for the X and 10 choices for the Y,
So that's 9*10 or 90, choice (C)


The next two palindromes greater than 50805 are m and  p  and  m < p . what is the value of  p – m
(A)  10    (B) 90   (C)  100  (D) 110 (E) 210

The next one is m=50905 and the next one after that is p=51015,
so p - m = 51015 - 50905 = 110, choice (D)

Edwin